OpenStudy (roadjester):

A tennis player player receives a shot with the ball (.06kg) travelling horizontally at 50.0 m/s and returns the shot with the ball travelling horizontally at 40.0 m/s in the opposite direction. a) What is the impulse delivered to the ball by the tennis racquet? b) What work does the racquet do on the ball?

OpenStudy (anonymous):

Eeek! I don't know...I'm not good with formulas and stuff, good luck though! :D

OpenStudy (apoorvk):

Impulse is change in momentum = final - initial So, initial momentum is = mv = 0.06 x 50 = 3 J m/sec final momentum = mv' = 0.06 x (-40) = -2.4J m/sec ...{the velocity is taken '-ve' due to change in direction. so impulse 'J' = 3 - (-2.4) = 5.4 J m/sec

OpenStudy (roadjester):

Is there any work done? I know that Work is the Force vector times displacement and the sum of forces is the derivative of momentum. the change in momentum/time is also the sum of forces.

OpenStudy (roadjester):

But what would be the displacement?

OpenStudy (apoorvk):

W.D. = change in kinetic energy = final - initial KE. = (1/2)mV^2

OpenStudy (roadjester):

Work-Kinetic Energy Theorem?

OpenStudy (apoorvk):

Exactly.

OpenStudy (roadjester):

What velocity do I use then?

OpenStudy (apoorvk):

you have both velocities stated - final and initial both.

OpenStudy (roadjester):

I know the squared will keep the answer positive.

OpenStudy (roadjester):

Ok...

OpenStudy (apoorvk):

Yes right.

OpenStudy (roadjester):

I think 40 because it's the return shot and that return shot is the one in which the racket strikes the ball.

OpenStudy (roadjester):

@apoorvk Would I use 40m/s as my velocity?

OpenStudy (apoorvk):

final will be 40, initial will be 50.

OpenStudy (apoorvk):

so find out initial and final kinetic energies, and then calculate the change.

OpenStudy (roadjester):

@apoorvk -27 Joules?

OpenStudy (apoorvk):

With Modulus!